OpenStudy (anonymous):

for the first half of a journey a car has a constant velocity of V1 and after that it has a constant velocity of V2. what is the average velocity?

OpenStudy (anonymous):

i was thinking V1+V2/2 but thats wrong :/

OpenStudy (anonymous):

i know that velocity= change in distance/change in time..

OpenStudy (anonymous):

This is my forte' ^^

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

so what do we do?

OpenStudy (anonymous):

let it travel with v1 for t1 covering s/2. & with v2 for t2 covering s/2. Okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So, v(avg) =s/t1+t2.

OpenStudy (anonymous):

Oops. silly mistake.. forgot the equations :0

OpenStudy (anonymous):

But did u see how i did it?

OpenStudy (anonymous):

write equations for t1 & t2 & add them up.

OpenStudy (anonymous):

okay brb

OpenStudy (anonymous):

@rajathsbhat where did your drawing go?

OpenStudy (anonymous):

I deleted it. It was wrong.

OpenStudy (anonymous):

v1=(s/2)/t1 then t1= (s/2)/v1 and v2=(s/2)/t2 then t2=(s/2)/v2 then average v: s/(s+(v1+v2) ? idk what am doing

OpenStudy (anonymous):